OpenStudy (anonymous):

If bill can clean the garage in four hours and john can clean the same garage in three hours, how many hours will it take to clean the garage if they work together?

OpenStudy (dmnts):

Garage - 1 Bills speed - 1/4h Johns speed - 1/3h Both speed: \[\frac{ 1 }{ 4 }+\frac{ 1 }{ 3 }=\frac{ 7 }{ 12 }=7/12h\] The answer is:\[\frac{ 12h }{ 7 }=1\frac{ 5 }{ 7 }h\]